OpenStudy (anonymous):

1. 4 - t = 3(t - 1) a. 3 b. 6 c. no solution d. identity 2. -6a + 3 = -3(2z - 1) a. a = 5 b. a = 10 c. no solution d. identity

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

You want to solve for t correct? on number 1 multiply the 3 with whats in parentheses 4-t=3t-3 then add 3 to both sides 7-t=3t add t to both sides 7=4t then divide by t 7/4 = t

OpenStudy (anonymous):

I know, I know, it's 1.75 = t but that's not on there. :(

OpenStudy (anonymous):

help? c:

OpenStudy (anonymous):

yea I'm leaning towards no solution bc identity means that the equation will equal itself with any value and putting in 3 gives 1=6 and putting in 6 gives -2=15

OpenStudy (anonymous):

It's funny cause I was thinking something similar to that! Ok! And the next on?

OpenStudy (anonymous):

yep looking at that now is that z suppose to be there or is it suppose to be an a

OpenStudy (anonymous):

it supposed to be a "a", my bad.

OpenStudy (anonymous):

its cool :) ok working on that now

OpenStudy (anonymous):

now this one is pretty easy to see one you foil out the -3

OpenStudy (anonymous):

\[-6a+3=-3(2a-1) \] \[-6a+3=-6a+3 \]

OpenStudy (anonymous):

Remember what I said early? Identity is an equation where you get the same results when you substitute any value. Meaning any number i put in this equation is going to be right. so plugging in 1 gives -3=-3 plugging in -2 is -9=-9 and so on and so forth. :)

OpenStudy (anonymous):

so it's a identity? Thanks !

OpenStudy (anonymous):

No problemo!

OpenStudy (anonymous):

If you need help with any others just leave me a message. I am pretty much always on here. :)

OpenStudy (anonymous):

hey, actually do you mind helping me on the last two ? then.

OpenStudy (anonymous):

Ask away!

OpenStudy (anonymous):

Thanks! ^^ give me a sec!

OpenStudy (anonymous):

3(c - 2) = 2(c - 6) a. 2 b. -6 c. no solution d. identity

OpenStudy (anonymous):

0.5(m + 4) = 3(m - 1) a. -2 b. 2 c. no solution d. identity

OpenStudy (anonymous):

That's all, I'm going to try to work them out of my side too! ^^

OpenStudy (anonymous):

ok well I have the answers but I want you to solve them yourself and post them step by step and i will tell you if its correct or where you went wrong.

OpenStudy (anonymous):

Wait. WHAT? noo help!! I'm going to go over all my homework tomorrow when I'm fully awoke soo pleassee!!!

OpenStudy (anonymous):

help/???? friieenndd??/

OpenStudy (anonymous):

okkaayy, give me a second!!!

OpenStudy (anonymous):

sorry was helping out another girl ok looking at the firs one 3c-6=2c-12 add 6 to both sides 3c=2c-6 subtract your 2c from both sides c=-6 on number 2 0.5m+2=3m-3 add 3 to both sides 0.5m+5=3m subtract your 0.5m from both sides 5=2.5m divide your 2.5 2=m

OpenStudy (anonymous):

..I was almost finish with it, you know. But thank you!!!! Thank you for all your help bro (or brolette!)!!!

OpenStudy (anonymous):

@cydneehill THANKS!!!

OpenStudy (anonymous):

It was no problem at all! I am always happy to help :)